factor completely and find the root of the expression
49x^2 -14x -3
Can you please help me with this?
Suzie

Found 2 solutions by Alan3354, htmentor:
Answer by Alan3354(69443)

Hint:
49 is either 1*49 or 7*7
3 is 1*3
------------

Answer by htmentor(1343)

factor completely and find the root of the expression
49x^2 -14x -3
=========================================
(7x + 1)(7x - 3)
So the roots are x = -1/7, x = 3/7